# Ravenea: A Benchmark for Multimodal Retrieval-Augmented Visual Culture Understanding

### v1

@@ 4.1 Culture-aware Contrastive Learning
Given an image $I_{i}$ associated with $T$ textual descriptions $\{D_{i1},D_{i2},\dots,D_{iT}\}$, each document $D_{it}$ is annotated with a binary label $y_{it}\in\{0,1\}$, where $y_{it}=1$ indicates cultural relevance and $y_{it}=0$ indicates irrelevance. For each image–text pair $(I_{i},D_{it})$, we employ a shared vision-language encoder—such as CLIP—to obtain modality-specific representations: $\mathbf{E}_{I_{i}}=\mathcal{E}_{V}(I_{i})$ for the visual input and $\mathbf{E}_{D_{it}}=\mathcal{E}_{L}(D_{it})$ for the textual input. We then compute the cosine similarity score $s_{it}$ between $\mathbf{E}_{I_{i}}$ and each corresponding $\mathbf{E}_{D_{it}}$, resulting in a similarity vector $\mathbf{S}_{i}=[s_{i1},s_{i2},\dots,s_{iT}]$.
Culture-awareness classification now amounts to:
$\mathcal{L}_{\text{Culture Classify}}=-\frac{1}{B\cdot T}\sum_{i=1}^{B}\sum_{j=1}^{T}\left[y_{ij}\log\sigma(s_{ij})+(1-y_{ij})\log(1-\sigma(s_{ij}))\right],$ (1)
where B is the number of images; $\sigma(\cdot)$ denotes the sigmoid function.
To prioritize culturally relevant descriptions in the ranking, we apply a margin ranking loss between all pairs of descriptions with differing cultural relevance. For each image $I_{i}$, we compare all pairs $(D_{ij},D_{ik})$ such that $y_{ij}=1$ and $y_{ik}=0$, and encourage the model to assign a higher similarity score to the relevant description. The ranking loss is defined as:
$\mathcal{L}_{\text{Rank}}=\frac{1}{B}\sum_{i=1}^{B}\sum_{\begin{subarray}{c}j,k=1\\ y_{ij}=1,y_{ik}=0\end{subarray}}^{T}\max\left(0,\delta-(s_{ij}-s_{ik})\right),$ (2)
@@
$\mathcal{L}_{\text{Diversity}}\left(\mathbf{S}_{i}\right)=-\sum_{t=1}^{T}\log\left(\frac{\exp(s_{it})}{\sum_{j=1}^{T}\exp(s_{ij})}\right)$ (3)
@@
Table 2: Performance with different retriever models. Fine-tuned contrastive models consistently outperform their frozen counterparts across tasks. Here, "CaSigLIP"=Culture-aware SigLIP2-SO/14@384px, and "CaCLIP"=Culture-aware CLIP-L/14@224px. Models in gray are frozen.
@@
Method MRR $\uparrow$ P@1 $\uparrow$ P@3 $\uparrow$ P@5 $\uparrow$ nDCG@1 $\uparrow$ nDCG@3 $\uparrow$ nDCG@5 $\uparrow$ SigLIP2-SO/14@384px 66.71 50.42 40.14 34.25 58.85 62.43 67.50 CLIP-L/14@224px 70.76 54.58 43.47 36.58 62.54 67.25 72.43 VisualBERT 59.66 42.50 35.97 32.75 51.29 55.49 62.29 VL-T5 55.53 35.42 34.03 31.42 45.21 53.18 59.62 LLaVA-OneVision-7B 54.15 36.20 31.83 28.69 45.85 50.82 56.74 CaSigLIP2 (ours) 69.35 54.17 43.47 36.50 61.98 66.75 71.98 CaCLIP (ours) 78.34 65.42 49.44 39.50 72.25 75.22 79.32
4.2 Multimodal Retrieval Results
@@
We perform a comprehensive evaluation of both frozen and fine-tuned retrievers, and present the results in Table 2. We find that fine-tuned models, particularly those based on contrastive learning, consistently outperform their frozen counterparts. For instance, CaCLIP achieves a substantial improvement in P@1, rising from 54.58% to 65.42%, and sets a new SOTA across all evaluation metrics. Although SigLIP2-SO/14@384px also benefits from fine-tuning, the performance gains are comparatively modest. In contrast, models such as LLaVA-OneVision-7B, VL-T5, and VisualBERT lag behind after fine-tuning, even underperforming relative to frozen baselines. This underperformance likely stems from the fact that models such as LLaVA-OneVision-7B and VisualBERT were originally pretrained for generative tasks with different objectives, whereas CLIP-L/14@224px and SigLIP2-SO/14@384px were explicitly trained for similarity-based alignment, providing them with a structural advantage in retrieval settings.
5 Multimodal Retrieval-augmented Visual Culture Understanding
We then evaluate the effectiveness of these retrievers with 14 SOTA VLMs, spanning a diverse set of architectures. We conduct experiments on two downstream tasks: cVQA and cIC, respectively.
5.1 Experimental Setup
Models.
We benchmark open and closed-weight SOTA VLMs on Ravenea, leveraging various retrievers against non-RAG baselines, assessing retrieval effectiveness across models of different sizes. The open-weight models include LLaVA-OneVision-7B li2025llavaonevision , Pixtral-12B agrawal2024pixtral , Phi-4 Multimodal-Instruct abouelenin2025phi , Gemma3-4B-Instruct and 27B-Instruct team2025gemma , Qwen2.5-VL-Instruct (3B, 7B, 72B  [^3] Due to computational constraints, we use the quantized version Qwen2.5-VL-Instruct-72B-AWQ.bai2025qwen2 , InternVL3 (2B, 8B, 38B) zhu2025internvl3 , and Deepseek-VL2 variants (Tiny and Base) wu2024deepseek . For the closed models, we adopt GPT-4.1 achiam2023gpt (accessed on 2025/04/14)  [^4] Knowledge cutoff: June 1, 2024; https://platform.openai.com/docs/models/gpt-4.1.
Evaluation metrics.
Table 3: Kendall’s $\tau$ rank correlation kendall1938new between automatic metrics and human judgments for the CCUB task. Statistically significant correlations ($p<0.05$) are marked with ✓. Our proposed metrics correlate stronger with human evaluation than the others.
Rouge-L lin-2004-rouge CIDER vedantam2015cider BERTScore Zhang*2020BERTScore: CLIPScore hessel-etal-2021-clipscore RegionScore (ours)
-0.172 ✗ -0.316 ✗ -0.011 ✗ 0.139 ✗ 0.442 ✓
For the cVQA task, we use accuracy as the primary evaluation metric, which measures the proportion of correctly predicted answers. For the cIC task, we employ several evaluation metrics including ROUGE-L lin-2004-rouge , CIDEr vedantam2015cider , BERTScore Zhang*2020BERTScore: , and CLIPScore hessel-etal-2021-clipscore , to assess the alignment between generated and reference captions across lexical, syntactic, and embedding-based levels. To further evaluate the cultural relevance and human-perceived quality, we conduct a human evaluation study. We employed four researchers to select the most accurate caption from 14 VLMs (see details in Appendix I and L), and find a significant mismatch between automatic metric scores and human judgments of cultural appropriateness (see Table 3). To bridge this gap, we further introduce RegionScore, a novel evaluation metric designed to quantify cultural grounding (see details in Appendix H and Table 9). It measures how well captions identify the correct country names tied to cultural elements, adding geographic and cultural specificity in image captioning.
5.2 Overall Performance
We present the main results in Table 4. The results demonstrate the efficacy of incorporating culture-aware retrieval augmentation. Employing fine-tuned retrievers yields substantial performance gains over both non-RAG and frozen retrievers baselines. Specifically, the CaCLIP achieves the highest average performance across both tasks, improves the accuracy for cVQA from 67.7% to 71.5%, and substantially improving the RegionScore for cIC from 40.2% to 58.1%. While CLIP-L/14@224px also offers improvements, fine-tuning consistently unlocks further potential. Furthermore, in the cVQA task, among all evaluated models, GPT-4.1 achieves the highest accuracy (86.8%) without RAG. Within the category of open-weight models, Qwen2.5-VL-72B leads with an accuracy of 81.0%. For lightweight models ($\leq$ 8B parameters), Qwen2.5-VL-7B achieves the best performance without RAG, reaching 67.7%. However, incorporating a CaCLIP significantly boosts performance, which enables InternVL3-8B to achieve 74.2% and outperform Qwen2.5-VL-7B by 0.6% with identical reranking. Notably, across multiple model families, augmenting the smallest variant with CaCLIP consistently elevates its performance to match or even exceed that of the next larger model tier. In the cIC task, with culture-aware contrastive learning, CaCLIP demonstrates substantial gains in identifying the culture in country-level of visual content, especially when built on top of VLMs with strong vision-language priors. It achieves the highest average RegionScore (58.1%) among the six reranking methods evaluated, with peak performance reaching 76.3% on the Gemma3-4B backbone. CaCLIP achieves leading the scores on 9 of 14 diverse suits of VLMs. This result underscores CaCLIP’s robustness and adaptability, particularly in culture-aware image captioning and retrieval tasks that demand fine-grained multimodal alignment.
@@ 5.2 Overall Performance
Table 4: cVQA and cIC Performance w/ and w/o RAG. Models in gray are frozen retrievers. Results are colored as Best. VLMs augmented with finetuned retriever generally perform better.
@@
Open Weights Closed Weights Retriever Average DeepSeek-VL2-Tiny DeepSeek-VL2 Qwen2.5-VL-3B Qwen2.5-VL-7B Qwen2.5-VL-72B InternVL3-2B InternVL3-8B InternVL3-38B Gemma3-4B Gemma3-27B Phi4-Multimodal Pixtral-12B LLaVA-OneVision-7B GPT-4.1 $cVQA_{Accuracy\uparrow}$ W/O RAG 67.7 57.1 65.5 64.2 67.7 81.0 58.4 66.5 71.9 63.6 78.7 55.5 66.1 64.8 86.8 SigLIP2-SO/14@384px 67.2 58.7 55.8 63.9 67.7 77.1 62.3 68.4 71.9 60.7 71.9 62.6 69.4 67.1 82.6 CLIP-L/14@224px 70.6 64.2 56.1 67.7 73.2 79.0 68.1 71.0 76.1 66.1 74.5 69.4 71.0 69.7 81.9 Finetuned Models VisualBERT 67.1 58.7 59.4 66.5 68.4 77.1 61.3 66.1 71.9 60.7 73.6 63.2 66.5 63.9 82.3 VL-T5 65.8 56.5 61.3 62.9 69.4 77.1 60.3 66.8 70.3 55.8 74.2 61.0 61.0 61.9 82.6 LLaVA-OneVision-7B 67.8 60.0 61.3 64.8 71.3 80.7 61.3 64.8 73.9 62.3 73.6 62.6 67.4 63.9 81.3 CaSigLIP2 (ours) 69.8 62.6 58.7 67.1 72.3 77.7 65.8 72.6 75.8 64.2 75.5 64.2 70.7 69.4 82.6 CaCLIP (ours) 71.5 65.8 60.0 69.4 73.6 81.0 68.7 74.2 75.2 66.8 75.2 66.8 71.0 70.0 83.2 $cIC_{RegionScore\uparrow}$ W/O RAG 40.2 31.3 36.3 36.3 48.8 47.5 27.5 37.5 22.5 67.5 62.5 10.0 35.0 18.8 47.5 SigLIP2-SO/14@384px 52.5 50.0 56.3 36.3 57.5 53.8 58.8 56.3 61.3 70.0 66.3 30.0 53.8 41.3 53.8 CLIP-L/14@224px 55.4 56.3 63.8 41.3 66.3 60.0 61.3 62.5 67.5 70.0 70.0 31.3 55.0 47.5 58.8 Finetuned Models VisualBERT 56.3 58.8 63.8 41.3 58.8 62.5 61.3 61.3 66.3 71.3 70.0 30.0 58.8 45.0 62.5 VL-T5 55.4 60.0 65.0 41.3 62.5 66.3 62.5 58.8 67.5 71.3 68.8 30.0 61.3 40.0 61.3 LLaVA-OneVision-7B 56.0 60.0 65.0 40.0 60.0 62.5 61.3 62.5 66.3 71.3 71.3 28.8 57.5 47.5 60.0 CaSigLIP2 (ours) 56.3 57.5 65.0 36.3 61.3 63.8 58.8 61.3 67.5 71.3 70.0 27.5 57.5 46.3 65.0 CaCLIP (ours) 58.1 60.0 67.5 42.5 65.0 65.0 62.5 60.0 70.0 76.3 75.0 27.5 62.5 52.5 55.0
6 Analysis and Further Discussion
@@
Culture-aware retrieval augmentation substantially benefits VLMs across both cVQA and cIC tasks, compared to their no-RAG counterparts. In this section, we explore the margin of the improvement, cultural preference and the effectiveness of cultural annotation.
6.1 Scaling Models Yields Diminishing or Negative Returns across Retrievers
@@
Figure 5: Performance improvements for smallest and largest models per family with multimodal retrievers. Scaling models yields marginal gains with various retrievers, even negative effects in both cVQA and cIC tasks. "ACC." denotes accuracy; "R.S." refers to the RegionScore; "$\Delta$" represents the change incorporated with RAG compared to the non-RAG baseline.
In the cVQA task, within the same VLM family, performance differences between RAG and non-RAG approaches exhibit non-monotonic trends as model size scales. For all four model families, larger models show marginal or even negative returns from RAG integration. What’s more, sensitivity to RAG varies across model families. Notably, DeepSeek-VL2 demonstrates the most pronounced performance gap: the smallest model benefits from RAG with an average improvement of approximately +5%, whereas the largest model in the same family suffers a degradation of around -6% on average. In the cIC task, the effectiveness of RAG exhibits a consistent trend with respect to model scale within a given model family. Across all four model families evaluated, larger models tend to benefit less—or at most comparably—from the integration of RAG, suggesting diminishing returns at higher capacity. Among them, Gemma3 models show the smallest relative improvement, achieving approximately a +7% performance gain on average, whereas InternVL3 models yield the highest benefit, with performance gains reaching up to +30%.
6.2 Differences across Countries
We evaluate all models using CaCLIP across a range of countries for both tasks, as shown in Figure 6. In the cVQA setting, most VLMs exhibit substantially diminished performance on culture-specific questions regarding Nigeria and Indonesia, in contrast to their performance on questions under other national contexts. Interestingly, questions related to Spanish culture reveal high inter-model variance,
Figure 6: Performance of 14 VLMs with CaCLIP across different countries. Despite the integration of CaCLIP, disparities in model performance persist across countries.
@@ 6.2 Differences across Countries
with accuracy differentials reaching up to 35%, underscoring significant discrepancies in cultural representation across models. In the cIC task, VLMs consistently underperform on images and documents associated with Indian cultural contexts, while achieving the highest RegionScores on Korean culture-related inputs. Model performance on Indian culture is particularly volatile, indicating inconsistent cultural grounding across architectures. By comparison, Korean and Chinese cultural inputs yield more stable performance across models, suggesting entrenched model-specific preferences in cultural alignment (see more results in Appendix J).
6.3 Ablation Study on Annotation Questions
Figure 7: Ablation for different annotation questions. Combining all three culture-relevant annotation questions yields the best performance.
@@ 6.3 Ablation Study on Annotation Questions
We further perform ablation studies across diverse combinations of annotation questions to assess their impact on downstream performance. Specifically, we evaluate 13 open-weight VLMs equipped with either CaSigLIP or CaCLIP, each trained on datasets constructed using varying subsets of culture-relevant annotations. From Figure 7, we can observe that leveraging all three questions (Q1 regarding country association; Q2 for topic alignment; Q3 for visual representation fidelity) yields the strongest performance on both cVQA and cIC tasks. For the cVQA task, we find Q1 provides the most significant benefit to CaSigLIP, whereas CaCLIP gains more from Q2. Among all pairwise combinations, the joint supervision from Q1 (country associations) and Q2 (topic alignment) proves slightly more effective than other pairs. In the cIC task, both CaSigLIP and CaCLIP achieve better performance improvements when trained with data derived from Q1, compared to other single-question sets. For pairwise combinations, CaCLIP benefits most from the Q1+Q3 combination, while CaSigLIP shows a clear preference for the Q2+Q3 setup.

### v2

@@ 3.2 Culture-aware Contrastive Learning
Given an image $I_{i}$ associated with a set of textual documents ${\mathbb{D}}=\{D_{i}^{1},D_{i}^{2},\dots,D_{i}^{n}\}$, each document $D_{i}^{j}$ is annotated with a binary label $y_{i}^{j}\in\{0,1\}$, where $y_{i}^{j}=1$ indicates cultural relevance and $y_{i}^{j}=0$ indicates irrelevance. For each image-document pair $(I_{i},D_{i}^{j})$, we employ a shared vision-language encoder, such as CLIP, to obtain modality-specific representations: $\mathbf{E}_{I_{i}}=\mathcal{E}_{V}(I_{i})$ for the visual input and $\mathbf{E}_{D_{i}^{j}}=\mathcal{E}_{L}(D_{i}^{j})$ for the textual input. We then compute the cosine similarity score $s_{i}^{t}$ between $\mathbf{E}_{I_{i}}$ and each corresponding $\mathbf{E}_{D_{i}^{j}}$, resulting in a similarity vector $\mathbf{S}_{i}=[s_{i}^{1},s_{i}^{2},\dots,s_{i}^{n}]$. Culture-awareness classification now amounts to:
$\mathcal{L}_{\text{Culture Classify}}=-\frac{1}{|{\mathbb{I}}|\cdot|{\mathbb{D}}|}\sum_{I_{i}\in{\mathbb{I}}}\sum_{D_{i}^{j}\in{\mathbb{D}}}\left[y_{i}^{j}\log\sigma(s_{i}^{j})+(1-y_{i}^{j})\log(1-\sigma(s_{i}^{j}))\right],$ (3)
where $\sigma(\cdot)$ denotes the sigmoid function.
To prioritize culturally relevant descriptions in the ranking, we apply a margin ranking loss between all pairs of descriptions with differing cultural relevance. For each image $I_{i}$, we compare all pairs $(D_{i}^{j},D_{i}^{k})$ such that $y_{i}^{j}=1$ and $y_{i}^{k}=0$, and encourage the model to assign a higher similarity score to the relevant description. The ranking loss is defined as:
$\mathcal{L}_{\text{Rank}}=\frac{1}{|{\mathbb{I}}|}\sum_{i=1}^{|{\mathbb{I}}|}\sum_{\begin{subarray}{c}j,k=1,\\ y_{i}^{j}=1,y_{i}^{k}=0\end{subarray}}^{|{\mathbb{D}}|}\max\left(0,\delta-(s_{i}^{j}-s_{i}^{k})\right),$ (4)
@@
$\mathcal{L}_{\text{Diversity}}\left(\mathbf{S}_{i}\right)=-\sum_{t=1}^{|{\mathbb{D}}|}\log\left(\frac{\exp(s_{i}^{t})}{\sum_{j=1}^{|{\mathbb{D}}|}\exp(s_{i}^{j})}\right)$ (5)
@@
Table 2: Performance with different retriever models. Fine-tuned contrastive models consistently outperform their frozen counterparts (shown in gray). Ravenea-SigLIP and Ravenea-CLIP are SigLIP2-SO/14@384px and CLIP-L/14@224px models fine-tuned on Ravenea, respectively.
@@
Method MRR $\uparrow$ P@1 $\uparrow$ P@3 $\uparrow$ P@5 $\uparrow$ nDCG@1 $\uparrow$ nDCG@3 $\uparrow$ nDCG@5 $\uparrow$ SigLIP2-SO/14@384px 68.62 54.66 37.47 32.92 61.22 63.82 71.44 CLIP-L/14@224px 75.44 60.87 41.41 34.41 67.75 72.31 78.09 VisualBERT 61.33 46.71 33.14 29.42 55.76 59.42 65.92 VL-T5 58.33 39.86 32.58 29.35 47.74 57.29 62.12 LLaVA-OV-7B 58.85 37.48 30.68 28.21 48.59 51.80 60.34 Ravenea-SigLIP (ours) 70.95 57.14 40.99 33.29 63.86 68.31 73.92 Ravenea-CLIP (ours) 82.17 72.05 45.76 36.77 77.08 78.96 84.09
4 Experimental Results
4.1 Main Results
@@
Multimodal retrieval results.   We perform a comprehensive evaluation of both frozen and fine-tuned retrievers, and present the results in Table 2. We find that fine-tuned models, particularly those based on contrastive learning, consistently outperform their frozen counterparts. For instance, Ravenea-CLIP achieves a substantial improvement in P@1, rising from 60.87% to 72.05%, and sets a new SOTA across all evaluation metrics. Although SigLIP2-SO/14@384px (Tschannen et al., 2025) also benefits from fine-tuning, the performance gains are comparatively modest. In contrast, models such as LLaVA-OV-7B (Li et al., 2025), VL-T5 (Cho et al., 2021), and VisualBERT (Li et al., 2019; 2020) lag behind after fine-tuning, even underperforming relative to frozen baselines. This underperformance likely stems from the fact that models such as LLaVA-OV-7B and VisualBERT were originally pretrained for generative tasks with different objectives, whereas CLIP-L/14@224px (Radford et al., 2021) and SigLIP2-SO/14@384px were explicitly trained for similarity-based alignment, providing them with a structural advantage in retrieval settings.
@@
Table 3: cVQA and cIC Performance w/ and w/o RAG. Models in gray are frozen retrievers. The best results are in bold. VLMs augmented with finetuned retriever generally perform better.
@@
Open-source Proprietary Retriever Average DeepSeek-VL2-Tiny DeepSeek-VL2 Qwen2.5-VL-3B Qwen2.5-VL-7B Qwen2.5-VL-72B InternVL3-2B InternVL3-8B InternVL3-78B Gemma3-4B Gemma3-27B Phi4-Multimodal Pixtral-12B LLaVA-OV-7B Qwen3-VL-2B Qwen3-VL-8B Qwen3-VL-32B GPT-4.1 $cVQA_{Accuracy\uparrow}$ W/O RAG 67.1 49.8 68.9 62.7 52.6 83.3 64.1 71.3 84.2 66.5 79.4 42.6 73.2 62.7 37.8 73.7 81.3 86.6 SigLIP2-SO/14@384px 71.0 48.8 76.1 63.2 67.5 82.3 71.3 73.7 86.1 67.5 77.5 64.6 75.1 44.5 67.9 75.6 83.3 82.8 CLIP-L/14@224px 71.9 50.2 77.5 63.6 66.0 82.8 73.7 78.5 84.2 69.4 76.6 67.0 76.6 44.5 68.9 76.6 83.3 83.3 Finetuned Models VisualBERT 67.6 48.8 71.8 61.7 64.1 78.5 66.5 67.5 82.3 63.6 75.1 61.2 68.4 37.3 60.8 73.2 83.7 85.2 VL-T5 66.1 46.6 72.7 58.4 65.1 78.5 65.1 67.9 80.4 58.4 75.6 58.9 62.7 35.4 60.3 75.6 78.9 83.7 LLaVA-OV-7B 67.9 50.1 66.0 60.3 67.0 81.8 66.5 66.0 84.2 65.1 75.1 60.8 68.9 37.3 59.8 77.0 83.7 84.7 Ravenea-SigLIP (ours) 72.6 48.8 76.1 68.4 67.9 81.3 73.2 75.6 85.6 70.3 79.4 69.4 75.6 47.8 68.9 77.0 84.2 85.6 Ravenea-CLIP (ours) 73.3 50.7 75.1 64.1 69.9 82.3 76.1 81.3 85.2 69.9 79.9 69.4 75.1 50.2 69.4 77.5 84.2 86.1 $cIC_{RegionScore\uparrow}$ W/O RAG 57.5 26.5 63.3 53.1 69.4 71.4 38.8 49.0 65.3 73.5 75.5 28.6 53.1 49.0 59.2 61.2 73.5 67.3 SigLIP2-SO/14@384px 54.8 30.6 59.2 51.0 63.3 65.3 44.9 53.1 57.1 65.3 69.4 46.9 53.1 42.9 49.0 59.2 59.2 61.2 CLIP-L/14@224px 65.1 38.8 69.4 59.2 71.4 69.4 65.3 65.3 71.4 73.5 73.5 61.2 69.4 51.0 63.3 67.3 71.4 65.3 Finetuned Models VisualBERT 57.4 36.7 57.1 53.1 61.2 69.4 44.9 59.2 61.2 63.3 69.4 42.9 57.1 44.9 55.1 65.3 69.4 65.3 VL-T5 57.0 34.7 61.2 55.1 65.3 73.5 46.9 53.1 61.2 61.2 67.3 44.9 59.2 38.8 53.1 63.3 67.3 63.3 LLaVA-OV-7B 56.8 38.8 63.3 51.0 63.3 69.4 44.9 57.1 59.2 65.3 71.4 40.8 55.1 46.9 49.0 61.2 67.3 61.2 Ravenea-SigLIP (ours) 60.6 30.6 73.5 57.1 67.3 71.4 44.9 59.2 69.4 67.3 69.4 57.1 61.2 40.8 53.1 65.3 71.4 71.4 Ravenea-CLIP (ours) 68.8 46.9 73.5 61.2 77.6 75.5 69.4 67.3 73.5 75.5 75.5 63.3 71.4 55.1 67.3 73.5 73.5 69.4
Downstream tasks.   We present downstream performance on both cVQA and cIC tasks in Table 3. The results demonstrate the efficacy of incorporating culture-aware retrieval augmentation. Our proposed fine-tuned retrievers, Ravenea-SigLIP and Ravenea-CLIP, consistently achieve the highest scores across both cVQA and cIC tasks, frequently outperforming both the non-RAG baselines and frozen retrievers (SigLIP2 / CLIP). Specifically, averaged over all VLMs, Ravenea-CLIP achieves the highest performance for both tasks, improving +6.2% in cVQA accuracy (67.1% $\rightarrow$ 73.3%), and +11.3% in cIC RegionScore (57.5% $\rightarrow$ 68.8%). While CLIP-L/14@224px also offers improvements, culture-aware contrastive learning consistently unlocks further potential. Furthermore, without retrieval, larger VLMs still achieve strong absolute performance on both tasks (e.g., InternVL3-78B at 84.2% on cVQA, and Gemma3-27B at 75.5% on cIC). Retrieval contents, however, disproportionately benefits small and mid size models. For lightweight models ($\leq$ 8B parameters), Ravenea-CLIP narrows the gap to the larger models. For example, Qwen3-VL-8B improves to 77.5% on cVQA, approaching the performance of the models an order of magnitude larger, highlighting the potential of culture-aware retrieval.
4.2 Further Findings
@@
Culture-aware retrieval augmentation substantially benefits VLMs across both cVQA and cIC tasks, compared to their non-RAG counterparts. In this section, we explore the margin of improvement, cultural preference, the effectiveness of cultural annotation, and the impact of context length.
Finding 1. While larger VLMs see limited gains from different retrievers, lightweight VLMs benefit substantially from culturally grounded retrieval contexts.
@@
Figure 4: Performance improvements per model family with Ravenea-CLIP. Scaling VLMs yields marginal gains on both cVQA and cIC tasks. $\Delta$" represents the change with Ravenea-CLIP relative to the non-RAG baseline.
To evaluate the scaling behavior of VLMs incorporating with Ravenea-CLIP, we select four model families spanning from 2B to 78B parameters and quantify the performance improvement $\Delta$ relative to non-RAG baselines. As shown in Figure 4, the impact of Ravenea-CLIP does not scale monotonically with model size: smaller models exhibit the most substantial gains, while larger models (27B to 78B) exhibit only modest improvements. For instance, Qwen3-VL-2B achieves +31.6% absolute improvement on cVQA and +8.1% on cIC, whereas Qwen3-VL-32B shows only +2.9% better on cVQA and no gain on cIC. This trend indicates that larger models may already internalize much of the relevant knowledge, especially the Wikipedia content retrieved by Ravenea-CLIP, resulting in redundancy and a plateau in performance. Overall, these results suggest that Ravenea-CLIP serves as an effective plug-and-play enhancement for lightweight models on both cVQA and cIC tasks, providing diminishing yet consistent gains as model size increases.
Finding 2. With Ravenea, VLMs show cultural disparities with and without RAG. Moreover, large inter-model variance highlights model-specific biases.
Figure 5: Geographic variability in VLM performance. While Ravenea-CLIP yields gains in general, there’s still variance in both baseline and relative improvement across geographic regions.
@@
For geographic variability, we evaluate all models using Ravenea-CLIP across a range of countries for both tasks, as shown in Figure 5. For cVQA, most VLMs exhibit substantially diminished performance on culture-specific questions regarding Nigeria and Indonesia, in contrast to their performance on questions under other national contexts. Interestingly, questions related to Spanish culture reveal high inter-model variance, with accuracy differentials reaching up to 50%, underscoring significant discrepancies in cultural representation across models. For cIC, VLMs consistently underperform on images and documents associated with Mexico cultural contexts, while achieving the highest RegionScores on Indian culture-related inputs. Model performance on Mexico culture is particularly volatile, indicating inconsistent cultural grounding across architectures. By comparison, Korean and Chinese cultural inputs yield more stable performance across models, suggesting entrenched model-specific preferences in cultural alignment (see more results in Appendix J).
Finding 3. Generally, models benefit most when trained with the full set of culture-relevant annotations. Culturally grounded retrieval substantially outperforms random or irrelevant contexts, underscoring both the necessity of culture-aware contexts. Furthermore, more or longer retrieved context length yield diminishing gains.
Figure 6: Ablations for annotations and retrieval content. Combining all three culture-relevant annotation questions consistently achieves the highest performance on both downstream tasks, validating the effectiveness of Ravenea setup.
@@
Annotation questions ablations.   We further perform ablation studies across diverse combinations of annotation questions to assess their impact on downstream performance. Specifically, we evaluate 16 open-weight VLMs equipped with either SigLIP2-SO/14@384px or CLIP-L/14@224px, each fine-tuned on datasets constructed using varying subsets of culture-relevant annotations. From Figure 6(a), we can observe that leveraging all three questions (Q1 regarding country association; Q2 for topic alignment; Q3 for explicit visual representation) yields the strongest performance on both cVQA and cIC tasks. For the cVQA task, we find Q1 provides the most significant benefit to Ravenea-SigLIP, whereas Ravenea-CLIP gains more from Q2. Among all pairwise combinations, the joint supervision from Q1 and Q2 proves slightly more effective than other pairs. In the cIC task, both Ravenea-SigLIP and Ravenea-CLIP achieve better performance improvements when trained with data derived from Q1, compared to other single-question sets. For pairwise combinations, Ravenea-CLIP benefits most from the Q1+Q3 combination, while Ravenea-SigLIP shows a clear preference for the Q2+Q3 setup.
Context relevance ablations.   Recent work (Roth et al., 2023; Ma et al., 2025) suggests that randomly augmented prompts may yield performance gains even in the absence of meaningful knowledge, raising concerns that improvements attributed to retrieval could in fact stem from other factors, such as superficial prompt lengthening and so on. To rigorously assess the effectiveness of culturally grounded retrieval, we design two control conditions: random character sequences, culturally unrelated documents. For random character sequences, we randomly generate synthetic strings in which each string consists of 100 to 300 substrings, and each substring contains 4 to 8 randomly sampled alphabetic characters, ensuring comparable token budgets without introducing interpretable content. For culturally unrelated documents, we randomly sample hundreds of Wikipedia articles from countries not represented in Ravenea, thereby maintaining language structure and topical coherence while explicitly breaking cultural relevance. Together, these conditions isolate whether gains arise from the related cultural context or merely from the presence of additional text. As shown in Figure 6(b), models augmented with Ravenea-CLIP consistently outperform other configurations, highlighting both the effectiveness of cultural retrieval and the robustness of the Ravenea evaluation protocol. More results are provided in Appendix J Table 11.
Figure 7: Context length ablations for cVQA task. With Ravenea-CLIP, adding longer or multiple documents yields diminishing returns, highlighting the importance of the context length.
Context length ablations.   We also extend our study with a controlled analysis of retrieval length, comparing three additional context granularities: paragraph-level, full top-1 document, and full top-3 documents. As shown in Figure 7, providing a compact, high-salience context (Top-1 256 tokens) consistently achieves the best or tied-best performance across most model scales. In contrast, expanding the context to full documents or multiple documents often leads to marginal gains or degradation. This trend is particularly evident for lightweight models, which appear more sensitive to the context length. Even for larger models, additional context yields only marginal improvements despite substantially increasing input length. These results indicate that simply retrieving more content does not necessarily translate into better performance, especially for the lightweight models.
5 Related Work
Retrieval augmentation for cultural understanding.   Retrieval augmentation has demonstrated significant efficacy for culture-related NLP tasks (Conia et al., 2024; Hu et al., 2024; Fung et al., 2023). Some works have explored retrieval-augmented approaches to cultural question answering. One line of work implemented retrieval using the World Values Survey to enable in-context learning for culturally contextualized questions (Seo et al., 2025; Yadav et al., 2025b), while others enhanced cultural question answering by retrieving information from web searches and knowledge bases (Lertvittayakumjorn et al., 2025). In cross-cultural applications, retrieval has also been shown effective for adapting recipes across cultural contexts (Hu et al., 2024). For multimodal understanding, however, the direction of culture-relevant retrieval remains underexplored.
Vision-language culture datasets.   Recent studies examine cultural understanding in VLMs through multicultural VQA (Liu et al., 2021; Nayak et al., 2024; Liu et al., 2025). Some researchers focus on culturally grounded visual categories such as cuisine, where models must distinguish dishes that are visually similar yet culturally distinct, highlighting fine-grained, geographically situated knowledge gaps (Winata et al., 2025; Li et al., 2024). Socioeconomic signals have also been explored, for example by studying how models infer or misinfer income level and living standards from visual cues (Nwatu et al., 2025). Others benchmark cultural entity recognition using Wikipedia-based prompts (Nikandrou et al., 2025), curate culture-specific image sets for value-based tasks (Yadav et al., 2025b) or concept-based image retrieval without explicitly considering cultural relevance (Bhatia et al., 2024). While prior efforts highlight the limitations of current VLM, they mainly assess the performance on retrieval and downstream tasks respectively, which are not suitable for probing how RAG shapes visual-cultural understanding. Our work specifically aims to fill this research gap.